sq root(x^2+5)=x+3
Thank you

Answer by stanbon(48564)

sq root(x^2+5)=x+3
--------------------
Square both sides to get:
x^2+5 = x^2+6x+9
6x = -4
x = -2/3
---------------
Check:
Can x = -2/3 ?
sqrt((4/9)+5) = -2/3 + 3
sqrt(49/9) = 7/3
7/3 = 7/3
-------------
Final Ans: x = -2/3
========================
Cheers,
Stan H.